MICO 1 st February 2010 Terry Hart (MOM). MICO 1 st February 2010 17 June 2015 Decay Solenoid DS will not be ready for Machine Physics. Matt Hills spent.
Published byModified over 5 years ago
Presentation on theme: "MICO 1 st February 2010 Terry Hart (MOM). MICO 1 st February 2010 17 June 2015 Decay Solenoid DS will not be ready for Machine Physics. Matt Hills spent."— Presentation transcript:
MICO 1 st February 2010 17 June 2015 Decay Solenoid DS will not be ready for Machine Physics. Matt Hills spent much of last week trying to clean helium in cold trap. Unfortunately this was unsuccessful due to probable leak contaminating helium with nitrogen. Linde to be contacted today to ascertain where to go from here. Large implications wrt what we will be able to achieve during machine physics. Some work to be delayed until User Run Target Target 2 commenced operation in R78 on Monday 25 th. As of Monday early afternoon, achieved 557,000 actuations. - Record for target run in R78! Plan is to run target to failure. - Mechanical fault - Excessive bearing wear
MICO 1 st February 2010 17 June 2015 MICE Hall Placed mylar sheet on Q6 girder box Modified covers over quadrupole cables Completed airline connection to decay solenoid valves Jan. 26 Health & Safety audit of MICE Hall went well Quote for hydraulic unit (beam stop back-up): £ 2.5K Compressor stands material arrived, work has commenced ELECTRICAL PS switch added to target and wired in situ – Ian Mullacrane DAQ Jean – Sebastian at RAL last week rewiring DAQ in control room Work complete but rewiring makes TOF calibration data necessary More details from Jean-Sebastien’s talk
MICO 1 st February 2010 17 June 2015 BLOC Rota Thursday, Jan. 28 th meeting established BLOC duties and rota until April 2010 BLOC will be on-call 24/7 to assist with any beamline/target issues during MICE operations Rehearsal Rehearsal on afternoon of Thursday 28 th January. The target operated for a few pulses. All magnets except DS (Q1 – Q9, D1 and D2) were ramped up to nominal currents. The beam stop tested All equipment functioned as expected, no problems to report.
MICO 1 st February 2010 17 June 2015 Machine Physics Four dates confirmed: Fri 5 th February16:00 – 00:00 Establish that DAQ and detectors work Sun 7 th February 08:30 – 00:00 Commission luminosity monitors Mon 8 th February16:00 – 00:00 Nothing has planned, left for contingency Wed 10 th February08:30 – 00:00 16h Activation study at 2V beamloss. TOF calibration probable (DAQ rewiring) Proton absorber studies (maybe, check with Marco)